John Albert of Johnny’s Farm is trying to forecast the number of bushels of apples he sells. He has data for the last 8 weeks of sales. The data are listed below.

1

2

3

4

5

6

7

8

70

73

65

62

80

74

69

78



Provide a forecast for Period 9 utilizing a 4-week moving average.

a.

77.90

b.

77.40

c.

75.25

d.

73.65
Answer #1

Answer: c. 75.25

Explanation:

  • Forecast of a period with 4 period moving average = average of last 4 periods demand
  • Forecast for week 9= sum of demand from 5 to 8 / 4 =(80+74+69+78)/4 = 75.25

Please ask, if you have any doubts through the comment section. Do rate the answer